### v1

@@ Appendix D Hyperparameter search and validation
We conduct a hyperparameter search using U.S. data from the period 1973 to 1994, employing weekly sampling and 22-day ahead returns, consistent with the setup for the main results in the paper. The period from 1973 to 1989 is utilized for training models with various hyperparameter values we are exploring. The period from 1990 to 1994 is used to evaluate the trained models and select the best hyperparameter values. Forecasts for the validation period are generated using the best hyperparameter values identified in the previous step. To maintain consistency with the train-test split used for the main results of the paper, we retrain the model annually during the validation period and use the most recent model for forecasting. Specifically, given a set of hyperparameter values, the model is initially trained on the 1973-1989 training sample and then rolled forward by one year four times, always using the model for the subsequent year’s forecasts. We utilize the full sample given the greater number of available observations compared to the liquid sample, which provides more stable results. Lengthening the validation sample would result in a shorter training sample, which is undesirable. The training sample should cover at least one severe market-wide negative shock in equities, which occurred in 1987. This approach suggests there is some potential room for improvement in finding better hyperparameters for the liquid sample and/or for international regions.
It is not feasible to perform a full hyperparameter search for all possible hyperparameters of the neural networks. Therefore, we divide the hyperparameter search into two sequential steps and focus only on a subset of hyperparameters. The rest of the hyperparameters are fixed and are based on the optimal hyperparameters found previously in the literature, e.g., Gu et al. (2020) or Tobek and Hronec (2021). Table D1 shows the fixed hyperparameters and their values.
Table D1: Fixed Hyperparameters. This table lists the hyperparameters that are fixed during the hyperparameter search. The number of epochs is set to $100*(A/n)$, where $n$ is the number of observations available for training and $A$ is an adjustment constant equal to $3{\small,}000{\small,}000$ for the full sample and $1{\small,}500{\small,}000$ for the liquid sample, derived from the average number of observations in individual samples over time.
@@
We start with a grid search for optimization algorithm hyperparameters, namely the learning rate (LR) and dropout rate (DR). We test values of $0.1$, $0.001$, and $0.0001$ for LR and 00, $0.2$, and $0.4$ for DR, evaluating the average quantile loss for the validation sample. We use a two-stage feedforward neural network with two hidden layers, each with $128$ neurons, as a starting point to perform the first round of hyperparameter search. The architecture was chosen to provide slightly higher capacity than neural networks used in Gu et al. (2020) or Tobek and Hronec (2021). Our evaluation criterion is the average quantile loss during the validation period from 1990 to 1994. Table D2 displays the results of the hyperparameter search.
@@
The lowest average quantile loss is achieved for the learning rate of $0.0003$ and dropout rate of $0.2$, although dropout rates of $0.1$ and $0.3$ are not significantly worse than $0.2$. Fixing the learning rate and dropout rate at the optimal values achieved in the first part of the hyperparameter search, we perform the second part of the hyperparameter search to find an optimal architecture for the two-stage feedforward neural network, where we try different numbers of neurons and layers in the individual stages of the network. Table D3 shows the results of the architecture hyperparameter search.
@@
A smaller or greater number of neurons than $8$ in the second stage leads to worse results. A smaller number of neurons is preferred to prevent potential overfitting and to speed up learning. The network already needs to be heavily regularized to achieve satisfactory performance; it, therefore, should not require more learning capacity by adding more neurons.
The L1 and L2 penalties were selected in isolation as it is not computationally possible to add them as another hyper-parameter in the search. The chosen values proved to be extremely robust and to work well across all the network architectures. The L1 penalty for the first layer of the first stage and L1/L2 penalty for the first layer of the second stage were selected based on inspection of estimated weights on the 1973-1994 training sample. The L1 penalty for the first layer of the first stage was selected using a one layer network and inspecting how many weights are non-negative. [^14] Note that one layer neural network is just a standard quantile linear regression with L1 penalty on its weights that is estimated using stochastic gradient descent. The goal was to induce some selection of input variables and to limit the effects of collinearity. The L1 penalty for the first stage is $0.0001$, and it forces weights of about $100-150$ of the input variables to zero. [^15] The L1 penalty for the first stage is changed to $0.00001$ for neural network specification, where raw returns are directly predicted from the first stage to reflect the smaller scale of the predicted variable.
The L1 and L2 penalties for the second stage were chosen solely to limit collinearity issues without any focus on variable selection. The penalty was chosen the smallest possible while still leading to no obvious collinearlity issues. [^16] The goal was to not have a mix of very large negative and positive weights, which are a symptom of the collinearity. The L1 and L2 penalties for the second stage are the same at $0.00001$. The L1 penalty for second stage is approximately $10$ times smaller than the L1 penalty for the first stage. The scale of the standardized stock returns, serving as a label for the first stage, is close to $1$. The scale of the raw stock returns, serving as a label for the second stage, is roughly $10$ times smaller. The L1 penalties for both stages are, therefore, consistent given the scale of the variables used as an input for the loss function.
D.0.1 Validation

### v2

@@ Appendix C Additional empirical results
Table C4: Quantiles and the cross-section of stock returns in the U.S.: 2hNN benchmark This table shows the average and volatility of monthly returns and annualised Sharpe ratios for decile long, short and long-short equal-weighted portfolios over the period 1995 to 2018. For each $\tau$-quantile forecast, the long portfolio is formed by taking the top 10% of stocks with the highest forecast each month. The quantile forecast are from neural network with two hidden layers (2hNN). The short portfolio is formed by taking the bottom 10% of stocks with the lowest forecast each month. The long-short portfolio is created by taking the difference between the long and short portfolios. Values in brackets are t-statistics adjusted for Newey-West standard errors with 12 lags. Results are obtained separately for the full and liquid samples for the US only.
Full Sample
Long Short Long-short
$\tau$ Mean Vol SR Mean Vol SR Mean Vol SR
0.01 1.16 (4.91) 2.61 1.54 0.26 (0.35) 12.87 0.07 0.89 (1.23) 12.33 0.25
0.05 1.22 (5.25) 2.52 1.68 -0.03 (-0.04) 10.39 -0.01 1.25 (1.73) 12.03 0.36
0.1 1.27 (5.52) 2.50 1.76 -0.22 (-0.30) 12.70 -0.06 1.50 (2.05) 11.81 0.44
0.2 1.37 (6.02) 2.52 1.88 -0.56 (-0.75) 12.12 -0.16 1.94 (2.60) 11.39 0.59
0.3 1.58 (6.80) 2.76 1.98 -0.96 (-1.25) 11.88 -0.28 2.53 (3.31) 10.82 0.81
0.4 2.01 (7.51) 3.55 1.96 -1.37 (-1.81) 11.04 -0.43 3.37 (4.49) 9.65 1.21
0.5 2.73 (7.11) 5.63 1.68 -1.66 (-2.27) 10.27 -0.56 4.39 (6.59) 6.98 2.18
0.6 3.11 (6.06) 7.98 1.35 -1.79 (-2.80) 8.61 -0.72 4.90 (9.48) 4.66 3.64
0.7 3.01 (4.81) 10.22 1.02 -0.74 (-2.14) 4.01 -0.64 3.75 (6.33) 7.51 1.73
0.8 2.51 (3.37) 12.08 0.72 0.58 (2.18) 2.51 0.80 1.93 (2.58) 11.14 0.60
0.9 1.90 (2.35) 13.16 0.50 0.81 (3.57) 2.38 1.18 1.09 (1.35) 12.59 0.30
0.95 1.53 (1.90) 13.25 0.40 0.89 (4.18) 2.47 1.25 0.64 (0.80) 12.32 0.18
0.99 1.23 (1.57) 12.91 0.33 0.95 (4.68) 2.74 1.20 0.28 (0.37) 12.12 0.08
Liquid Sample
Long Short Long-short
$\tau$ Mean Vol SR Mean Vol SR Mean Vol SR
0.01 0.95 (5.25) 2.96 1.11 0.19 (0.25) 13.16 0.05 0.76 (1.03) 11.45 0.23
0.05 0.95 (5.37) 2.91 1.13 0.12 (0.15) 13.86 0.03 0.84 (1.14) 11.64 0.25
0.1 0.99 (5.60) 2.96 1.16 0.10 (0.14) 11.55 0.03 0.88 (1.20) 11.29 0.27
0.2 1.10 (6.18) 3.05 1.25 0.04 (0.06) 13.86 0.01 1.06 (1.48) 11.13 0.33
0.3 1.25 (6.49) 3.36 1.29 -0.01 (-0.02) 0.00 -0.00 1.26 (1.85) 10.39 0.42
0.4 1.42 (5.97) 4.10 1.20 -0.06 (-0.08) 10.39 -0.02 1.47 (2.39) 8.93 0.57
0.5 1.68 (4.81) 6.40 0.91 -0.10 (-0.17) 8.66 -0.04 1.78 (4.59) 6.42 0.96
0.6 1.54 (3.43) 8.60 0.62 0.11 (0.29) 6.35 0.06 1.42 (5.92) 6.65 0.74
0.7 1.24 (2.21) 10.23 0.42 0.46 (2.29) 3.39 0.47 0.78 (1.63) 9.01 0.30
0.8 1.03 (1.54) 11.51 0.31 0.65 (3.41) 2.92 0.77 0.38 (0.60) 10.97 0.12
0.9 0.77 (1.07) 12.12 0.22 0.74 (3.98) 2.88 0.89 0.03 (0.04) 10.39 0.01
0.95 0.61 (0.81) 12.43 0.17 0.75 (3.94) 2.89 0.90 -0.14 (-0.19) 12.12 -0.04
0.99 0.50 (0.67) 12.37 0.14 0.80 (4.23) 2.92 0.95 -0.30 (-0.41) 11.55 -0.09
Table C5: Moments decile portfolios: 2hNN benchmark. This table shows average monthly returns for individual deciles (1-10) and long-short equal-weighted portfolios built at the predicted moments. The quantile forecast are from neural network with two hidden layers (2hNN). Results are reported for the full and liquid samples, as well as for individual regions, covering the period from 1995 to 2018. Values in parentheses are t-statistics adjusted for Newey-West standard errors with 12 lags.
Mean Full Sample Liquid Sample
Decile USA Europe Japan Asia Pacific USA Europe Japan Asia Pacific
1 -1.98 (-3.18) -0.52 (-0.88) -1.07 (-1.64) -1.19 (-1.57) -0.01 (-0.01) -0.41 (-0.77) -0.77 (-1.45) -0.67 (-1.07)
2 -0.13 (-0.29) -0.16 (-0.33) -0.18 (-0.35) -0.05 (-0.08) 0.52 (1.41) 0.27 (0.66) -0.28 (-0.61) 0.26 (0.58)
3 0.51 (1.39) 0.07 (0.15) -0.01 (-0.03) 0.39 (0.70) 0.54 (1.73) 0.43 (1.09) -0.14 (-0.32) 0.47 (1.10)
4 0.76 (2.40) 0.36 (0.95) 0.17 (0.41) 0.74 (1.39) 0.74 (2.49) 0.63 (1.71) 0.10 (0.25) 0.72 (1.58)
5 0.93 (3.18) 0.67 (1.92) 0.38 (0.94) 0.92 (1.91) 0.95 (3.44) 0.78 (2.28) 0.31 (0.80) 0.81 (1.86)
6 1.23 (4.32) 0.95 (2.80) 0.60 (1.61) 1.36 (2.95) 0.97 (3.55) 1.03 (2.94) 0.33 (0.87) 1.02 (2.45)
7 1.38 (4.51) 1.18 (3.38) 0.89 (2.29) 1.78 (3.51) 1.11 (3.77) 0.97 (2.71) 0.57 (1.50) 1.09 (2.37)
8 1.73 (5.20) 1.47 (4.09) 1.16 (2.96) 2.23 (4.13) 1.06 (3.25) 1.06 (2.93) 0.62 (1.75) 1.01 (1.99)
9 2.08 (5.77) 1.85 (4.96) 1.58 (3.57) 2.74 (4.60) 1.29 (3.60) 1.26 (3.22) 0.74 (1.94) 1.06 (1.97)
10 3.60 (6.70) 3.65 (7.86) 2.37 (4.94) 5.24 (7.26) 1.66 (4.00) 1.45 (3.23) 0.92 (2.08) 1.54 (2.62)
10-1 5.59 (10.48) 4.17 (10.22) 3.44 (9.24) 6.43 (12.91) 1.67 (5.79) 1.86 (7.77) 1.68 (5.78) 2.22 (5.83)
Volatility
1 1.02 (4.71) 0.74 (3.06) 0.22 (0.85) 0.77 (2.58) 0.87 (4.77) 0.80 (3.46) 0.31 (1.41) 0.94 (3.01)
2 1.07 (4.12) 0.85 (2.95) 0.34 (1.06) 0.98 (2.49) 0.93 (4.00) 0.84 (2.92) 0.24 (0.77) 0.91 (2.51)
3 1.06 (3.93) 0.86 (2.74) 0.46 (1.33) 1.07 (2.38) 1.08 (4.39) 0.78 (2.47) 0.28 (0.85) 1.03 (2.47)
4 1.06 (3.66) 0.89 (2.54) 0.57 (1.53) 0.98 (1.99) 1.05 (4.02) 0.84 (2.49) 0.35 (1.00) 1.00 (2.39)
5 1.11 (3.47) 0.81 (2.17) 0.64 (1.53) 0.93 (1.68) 1.01 (3.49) 0.80 (2.24) 0.37 (1.04) 0.77 (1.70)
6 1.14 (3.11) 0.76 (1.86) 0.67 (1.51) 1.01 (1.72) 0.98 (3.37) 0.81 (2.09) 0.35 (0.90) 0.89 (1.77)
7 1.07 (2.46) 0.56 (1.24) 0.74 (1.49) 1.17 (1.76) 0.94 (2.53) 0.74 (1.89) 0.32 (0.73) 0.78 (1.42)
8 0.85 (1.64) 0.35 (0.70) 0.77 (1.39) 1.14 (1.57) 0.85 (1.86) 0.90 (1.94) 0.35 (0.72) 0.73 (1.31)
9 0.80 (1.25) 0.70 (1.18) 0.95 (1.52) 1.92 (2.30) 0.72 (1.27) 0.70 (1.28) 0.09 (0.17) 0.33 (0.51)
10 0.93 (1.19) 3.02 (4.00) 0.52 (0.74) 4.18 (4.70) 0.41 (0.54) 0.25 (0.36) -0.25 (-0.33) -0.07 (-0.08)
10-1 -0.09 (-0.12) 2.28 (3.16) 0.30 (0.53) 3.41 (4.64) -0.45 (-0.61) -0.55 (-0.93) -0.56 (-0.88) -1.00 (-1.44)
Skewness
1 1.00 (3.70) 0.77 (2.17) 0.35 (1.01) 0.73 (1.70) 0.84 (3.41) 0.69 (2.15) 0.24 (0.68) 0.80 (2.37)
2 1.05 (3.79) 0.77 (1.99) 0.36 (0.92) 0.83 (1.74) 0.92 (3.58) 0.76 (2.27) 0.24 (0.71) 0.86 (2.21)
3 1.02 (3.45) 0.68 (1.68) 0.40 (0.98) 0.82 (1.62) 1.02 (4.14) 0.81 (2.48) 0.35 (1.12) 0.75 (1.75)
4 1.07 (3.36) 0.58 (1.44) 0.45 (1.01) 0.83 (1.60) 1.05 (4.03) 0.75 (2.28) 0.33 (0.87) 0.80 (1.78)
5 1.02 (3.12) 0.56 (1.39) 0.59 (1.24) 0.98 (1.85) 1.09 (3.98) 0.83 (2.25) 0.38 (0.99) 0.56 (1.24)
6 0.98 (2.84) 0.61 (1.58) 0.64 (1.31) 1.46 (2.48) 1.01 (3.20) 0.81 (2.16) 0.22 (0.54) 0.82 (1.87)
7 0.97 (2.54) 0.78 (2.00) 0.78 (1.63) 1.65 (2.72) 0.88 (2.56) 0.75 (1.86) 0.37 (0.81) 0.71 (1.40)
8 1.19 (2.66) 1.04 (2.71) 0.88 (1.83) 2.27 (3.34) 0.74 (1.83) 0.85 (1.94) 0.36 (0.79) 0.70 (1.23)
9 1.10 (2.05) 1.63 (3.85) 0.92 (1.85) 2.77 (3.60) 0.71 (1.39) 0.71 (1.57) 0.04 (0.08) 0.70 (1.11)
10 0.69 (1.11) 2.10 (4.13) 0.52 (1.01) 1.82 (2.24) 0.59 (0.86) 0.51 (0.82) -0.13 (-0.22) 0.60 (0.81)
10-1 -0.30 (-0.62) 1.33 (3.66) 0.17 (0.54) 1.09 (2.00) -0.25 (-0.42) -0.18 (-0.43) -0.37 (-0.98) -0.20 (-0.37)
Kurtosis
1 1.06 (3.08) 0.71 (1.56) 0.33 (0.70) 0.61 (1.14) 1.15 (3.44) 0.89 (2.09) 0.51 (1.17) 0.86 (1.74)
2 1.04 (3.23) 0.77 (1.92) 0.43 (1.07) 0.73 (1.48) 1.04 (3.07) 0.87 (2.09) 0.31 (0.71) 0.73 (1.60)
3 1.02 (3.28) 0.75 (1.89) 0.44 (1.12) 0.85 (1.74) 0.90 (2.66) 0.79 (2.08) 0.32 (0.79) 0.81 (1.58)
4 1.09 (3.62) 0.66 (1.61) 0.48 (1.13) 0.92 (1.78) 1.01 (3.27) 0.76 (2.01) 0.30 (0.73) 0.78 (1.57)
5 1.04 (3.23) 0.71 (1.77) 0.66 (1.46) 1.03 (1.84) 0.98 (3.27) 0.65 (1.72) 0.25 (0.64) 0.63 (1.28)
6 0.96 (2.79) 0.69 (1.72) 0.70 (1.49) 1.62 (2.72) 0.88 (2.90) 0.72 (1.90) 0.27 (0.67) 0.76 (1.66)
7 0.98 (2.53) 0.81 (2.14) 0.85 (1.79) 1.88 (3.04) 0.80 (2.34) 0.82 (2.24) 0.23 (0.58) 0.75 (1.68)
8 1.07 (2.41) 1.09 (2.83) 0.86 (1.80) 2.45 (3.43) 0.74 (2.13) 0.83 (2.18) 0.15 (0.37) 0.86 (1.74)
9 1.05 (2.11) 1.48 (3.64) 0.77 (1.60) 2.46 (3.51) 0.69 (1.63) 0.59 (1.46) 0.20 (0.48) 0.63 (1.24)
10 0.79 (1.52) 1.87 (4.57) 0.35 (0.76) 1.59 (2.35) 0.64 (1.35) 0.54 (1.29) -0.13 (-0.30) 0.49 (0.83)
10-1 -0.27 (-0.75) 1.16 (3.80) 0.01 (0.05) 0.98 (2.29) -0.51 (-1.28) -0.35 (-1.50) -0.63 (-2.40) -0.37 (-0.93)
Table C6: Mean forecast evaluation: 2hNN benchmark. This table shows the $R^{2}$ of the mean and median forecasts of the next 22 days of stock returns for the full and liquid samples across all regions. The mean prediction based on the two-stage quantile model (Two-stage NN) is compared with the direct mean prediction from the two-hidden-layer neural network with MSE loss (2hNN MSE - Mean), derived mean predictions from the two-hidden-layer neural network with multi-head quantile loss (2hNN Quantile - Mean), and the median prediction from the two-stage model (Two-stage NN - Median). Results are obtained separately for full and liquid samples on individual regions covering the period from 1995 to 2018. Models are retrained annually using the expanding window for US data only, while out-of-sample forecasts are generated for both US and international data. Note that losses are multiplied by 100, and values in parentheses are Diebold-Mariano test statistics for the difference between the model in a row and the two-stage NN mean. Standard errors are adjusted using the Newey-West standard deviation estimator with 12 lags.
Specification Full Sample Liquid Sample
2hNN MSE - Mean 0.65 0.14 0.85 0.69 0.51 -0.68 -0.44 -0.48 -0.77 -0.54
2hNN Quantile - Mean 1.00 0.45 1.21 1.01 0.79 -0.98 -0.77 -0.89 -0.93 -0.92
(-1.95) (-2.54) (-1.53) (-3.02) (-2.68) (-1.86) (-1.78) (-2.24) (-1.86) (-2.21)
Two stage NN - Median 0.10 -0.04 0.30 0.06 0.09 0.37 0.32 0.47 0.55 0.50
Two stage NN - Mean 1.63 0.87 1.66 1.63 1.36 0.32 0.22 0.18 0.40 0.37
Table C7: Volatility forecast evaluation: 2hNN benchmark. This table shows the mean absolute deviation (MAD), root mean squared error (RMSE), and Diebold Mariano test statistic (DM) comparing the corresponding losses of the three competing models for the 22-day volatility forecasts. Volatility forecasting based on two-stage quantile forecasting (two-stage NN) is compared with GARCH model forecasting and volatility derived from quantile forecast from neural network with two hidden layers (2hNN). Results are obtained separately for full and liquid samples on individual regions covering the period from 1995 to 2018. Neural network models are retrained annually using an expanding window on US data only, while out-of-sample forecasts are generated for both US and international samples. Diebold-Mariano test statistics are computed for the difference between the two-stage NN model volatility forecasts and the two competing models - GARCH and 2hNN. Note that losses are multiplied by 100 and the standard errors used in the Diebold-Mariano test statistics are adjusted using the Newey-West standard deviation estimator with 12 lags.
Specification Variable Full Sample Liquid Sample
GARCH MAD 7.34 8.79 7.79 11.1 8.53 3.89 3.54 3.80 4.87 3.93
2hNN MAD 4.66 5.32 3.91 6.41 5.01 3.04 2.76 2.85 3.18 2.95
Two stage NN MAD 4.63 5.20 3.99 6.49 4.99 2.93 2.63 2.81 3.16 2.87
GARCH RMSE 13.14 16.24 14.63 18.70 15.50 6.93 7.10 6.72 9.52 7.36
2hNN RMSE 7.57 8.48 5.58 10.43 8.08 4.73 4.12 4.05 4.95 4.48
Two stage NN RMSE 7.46 8.52 5.64 10.05 8.08 4.62 4.03 3.98 4.86 4.38
GARCH DM (10.25) (35.92) (20.32) (44.26) (26.47) (7.90) (5.98) (7.97) (9.40) (10.02)
2hNN DM (1.56) (-1.12) (-1.69) (-1.03) (-0.18) (2.38) (1.59) (1.60) (1.41) (2.10)
@@ Appendix D Hyperparameter search and validation
We conduct a hyperparameter search using U.S. data from the period 1973 to 1994, employing weekly sampling and 22-day ahead returns, consistent with the setup for the main results in the paper. The period from 1973 to 1989 is utilized for training models with various hyperparameter values we are exploring. The period from 1990 to 1994 is used to evaluate the trained models and select the best hyperparameter values. Forecasts for the validation period are generated using the best hyperparameter values identified in the previous step. Hyperparameters are optimized once on the early period and held fixed throughout the evaluation horizon to limit the computational burden associated with repeated tuning. To maintain consistency with the train-test split used for the main results of the paper, we retrain the model annually during the validation period and use the most recent model for forecasting. Specifically, given a set of hyperparameter values, the model is initially trained on the 1973-1989 training sample and then rolled forward by one year four times, always using the model for the subsequent year’s forecasts. We utilize the full sample given the greater number of available observations compared to the liquid sample, which provides more stable results. Lengthening the validation sample would result in a shorter training sample, which is undesirable. The training sample should cover at least one severe market-wide negative shock in equities, which occurred in 1987. This approach suggests there is some potential room for improvement in finding better hyperparameters for the liquid sample and/or for international regions.
It is not feasible to perform a full hyperparameter search for all possible hyperparameters of the neural networks. Therefore, we divide the hyperparameter search into two sequential steps and focus only on a subset of hyperparameters. The rest of the hyperparameters are fixed and are based on the optimal hyperparameters found previously in the literature, e.g., Gu et al. (2020a) or Tobek and Hronec (2021). Table D1 shows the fixed hyperparameters and their values.
[…]
@@
We start with a grid search for optimization algorithm hyperparameters, namely the learning rate (LR) and dropout rate (DR). We test values of 0.10.10.1, 0.0010.0010.001, and 0.00010.00010.0001 for LR and 0, 0.20.20.2, and 0.40.40.4 for DR, evaluating the average quantile loss for the validation sample. We use a two-stage feedforward neural network with two hidden layers, each with 128128128 neurons, as a starting point to perform the first round of hyperparameter search. The architecture was chosen to provide slightly higher capacity than neural networks used in Gu et al. (2020a) or Tobek and Hronec (2021). Our evaluation criterion is the average quantile loss during the validation period from 1990 to 1994. Table D2 displays the results of the hyperparameter search.
@@
The lowest average quantile loss is achieved for the learning rate of 0.00030.00030.0003 and dropout rate of 0.20.20.2, although dropout rates of 0.10.10.1 and 0.30.30.3 are not significantly worse than 0.20.20.2. Fixing the learning rate and dropout rate at the optimal values achieved in the first part of the hyperparameter search, we perform the second part of the hyperparameter search to find an optimal architecture for the two-stage feedforward neural network, where we try different numbers of neurons and layers in the individual stages of the network. Table D3 shows the results of the architecture hyperparameter search.
@@
A smaller or greater number of neurons than 888 in the second stage leads to worse results. A smaller number of neurons is preferred to prevent potential overfitting and to speed up learning. The network already needs to be heavily regularized to achieve satisfactory performance; it, therefore, should not require more learning capacity by adding more neurons.
The L1 and L2 penalties were selected in isolation as it is not computationally possible to add them as another hyper-parameter in the search. The chosen values proved to be extremely robust and to work well across all the network architectures. The L1 penalty for the first layer of the first stage and L1/L2 penalty for the first layer of the second stage were selected based on inspection of estimated weights on the 1973-1994 training sample. The L1 penalty for the first layer of the first stage was selected using a one layer network and inspecting how many weights are non-negative. [^16] Note that one layer neural network is just a standard quantile linear regression with L1 penalty on its weights that is estimated using stochastic gradient descent. The goal was to induce some selection of input variables and to limit the effects of collinearity. The L1 penalty for the first stage is 0.00010.00010.0001, and it forces weights of about 100150100-150100 - 150 of the input variables to zero. [^17] The L1 penalty for the first stage is changed to 0.000010.000010.00001 for neural network specification, where raw returns are directly predicted from the first stage to reflect the smaller scale of the predicted variable.
The L1 and L2 penalties for the second stage were chosen solely to limit collinearity issues without any focus on variable selection. The penalty was chosen the smallest possible while still leading to no obvious collinearlity issues. [^18] The goal was to not have a mix of very large negative and positive weights, which are a symptom of the collinearity. The L1 and L2 penalties for the second stage are the same at 0.000010.000010.00001. The L1 penalty for second stage is approximately 101010 times smaller than the L1 penalty for the first stage. The scale of the standardised stock returns, serving as a label for the first stage, is close to 111. The scale of the raw stock returns, serving as a label for the second stage, is roughly 101010 times smaller. The L1 penalties for both stages are, therefore, consistent given the scale of the variables used as an input for the loss function.
D.1 Validation